# GMSR:Gradient-Guided Mamba for Spectral Reconstruction from RGB Images

### v1

@@ Abstract
Mainstream approaches to spectral reconstruction (SR) primarily focus on designing Convolution- and Transformer-based architectures. However, CNN methods often face challenges in handling long-range dependencies, whereas Transformers are constrained by computational efficiency limitations. Recent breakthroughs in state-space model (e.g., Mamba) has attracted significant attention due to its near-linear computational efficiency and superior performance, prompting our investigation into its potential for SR problem. To this end, we propose the Gradient-guided Mamba for Spectral Reconstruction from RGB Images, dubbed GMSR-Net. GMSR-Net is a lightweight model characterized by a global receptive field and linear computational complexity. Its core comprises multiple stacked Gradient Mamba (GM) blocks, each featuring a tri-branch structure. In addition to benefiting from efficient global feature representation by Mamba block, we further innovatively introduce spatial gradient attention and spectral gradient attention to guide the reconstruction of spatial and spectral cues. GMSR-Net demonstrates a significant accuracy-efficiency trade-off, achieving state-of-the-art performance while markedly reducing the number of parameters and computational burdens. Compared to existing approaches, GMSR-Net slashes parameters and FLOPS by substantial margins of 10 times and 20 times, respectively. Code is available at https://github.com/wxy11-27/GMSR.
@@
Lightweight Spectral Reconstruction, Mamba, State Space Model.
@@ I Introduction
Hyperspectral imaging is a technique for acquiring images by capturing spectral information of an object at multiple contiguous wavelengths. Hyperspectral images can provide spectral information beyond the visible range of the human eye, thus making it possible to analyze and identify the properties, composition, and state of objects in greater detail. The fine spectral information of HSIs can significantly improve the accuracy of feature classification and identification and is widely used for target identification [1], environmental monitoring [2], and resource investigation [3]. However, hyperspectral imaging devices are often expensive, leading to high data acquisition costs, which limits their widespread use in certain fields. Additionally, due to the inherent physical properties of hyperspectral imaging sensors, hyperspectral images (HSIs) always suffer from complex noise, low signal-to-noise ratio, and low spatial resolution, greatly restricting the accuracy improvement of high-level remote sensing tasks.
Snapshot Compressive Imaging (SCI) systems and computational reconstruction algorithms [4, 5, 6] are effective means for reconstructing three-dimensional HSI cubes. However, these methods still rely on expensive hardware equipment. As a cost-effective alternative, spectral super-resolution (SSR) or spectral reconstruction (SR) techniques have attracted attention. SR techniques use color information and known spectral features in RGB images to estimate the spectral response curve of each pixel point to generate HSIs. The current SR methods can roughly be divided into sparse dictionary-learning [7, 8] and deep learning-based methods [9, 10].
@@
Figure 1: PSNR-Params-FLOPS comparisons with other spectral reconstruction algorithms. The FLOPS (computational complexity) is represented on the horizontal axis, while the vertical axis represents PSNR (performance), with a circle radius indicating Params (memory cost). Our Gradient-Guided Mamba (GMSR-Net) outperforms other methods with lower FLOPS and Params requirements.
@@
The early SR research primarily focused on establishing sparse coding or relatively shallow learning models within specific spectral bands. However, due to their limited expressive power, these methods could only be effective within constrained image domains. In recent years, advancements in deep learning have significantly propelled the development of SR, overcoming many limitations and achieving significant progress. Convolutional neural networks (CNNs) have become the mainstream solution for SR due to their powerful representation and fitting capabilities. Current CNN-based methods treat SR as a regression problem, achieving it by establishing a mapping between HSIs and RGB images. Despite achieving significant performance improvements, CNN-based methods have limitations in capturing long-range dependencies and spectral self-similarity. The Transformer [11] model was originally designed for natural language processing (NLP) tasks, but has since been successfully applied to computer vision. Its advantage lies in its ability to handle long-range dependencies and focus on global information. Cai et al.  [5] first proposed the end-to-end framework MST based on Transformer, which aims to reconstruct HSI through compressed measurements. Building on this, Lin et al. [12] embedded the sparsity of HSI into the Transformer model and developed a fine learning scheme for spectral synthesis imaging. However, since the Transformer needs to compute the attention mechanism between each image patch, the complexity of computation increases exponentially as the resolution increases. Constrained by computational resources, it is usually necessary to divide the image into smaller patches. This preprocessing step inevitably results in each patch contains only a portion of the object, thus providing limited contextual information, which damages the accuracy of reconstruction.
@@
Recently, a novel architecture named Mamba has attracted widespread attention. Compared to traditional Transformers, Mamba demonstrates greater efficiency in handling long sequences while maintaining training and inference effectiveness. Scholars have begun to introduce Mamba into the field of remote sensing image processing, achieving innovative progress in tasks such as image classification [13] and semantic segmentation [14]. However, the aforementioned tasks emphasize high-level information in images, such as objects, scenes, and semantic labels. The capability of Mamba to reconstruct low-level information such as colors, textures, and edges is still underexplored and calls for further investigation. This also inspires us to delve deeper into the potential of Mamba in remote modeling within HSI spectral reconstruction networks.
@@
Mamba requires scanning the image row-by-row or column-by-column, which leads to the separation of tokens that are otherwise adjacent in 2D space, thus increasing the difficulty of feature extraction in localized regions. However, in HSI, local details such as edges and textures are crucial for subsequent analysis and understanding. The gradient map is a great means to effectively reflect the edge and texture information in an image. Currently, the extraction and utilization of image gradients has been widely recognized as a key technique in several fields such as target detection [15] and image segmentation [16]. For example, Canny [15] has become a classical edge detection method based on gradient information by utilizing the gradient information of an image at multiple scales. The SPSR [17] employs an additional gradient branching network to generate a high-resolution gradient map as an additional structural prior, thus alleviating the geometric deformation problem. Distinguished from natural images, HSIs contain not only spatial structural information but also rich spectral information. The gradient map in the spectral dimension reflects the variation between different spectral channels in HSI. Currently, the utilization of gradient information in the spectral dimension of HSI remains relatively limited, and the extraction of spectral information still needs to be further improved.
@@
To overcome the above challenges, we introduce the state space model(SSM) into spectral reconstruction tasks for the first time and propose GMSR-Net. GMSR-Net is an efficient spectral reconstruction model with a global receptive field and linear complexity, consisting of multiple Gradient-Guided Mamba (GM) blocks. Each GM block utilizes a three-branch structure. In addition to efficient global feature extraction using Mamba, spatial structural attention and spectral structural attention are innovatively introduced to fully maintain structural information in both spatial and spectral dimensions. It is worth noting that GMSR-Net successfully balances accuracy and computational efficiency. As shown in Fig. 1, it achieves state-of-the-art performance while significantly reducing parameters and computational costs. The main contributions of this paper are:
@@
  • We propose GMSR-Net, the first Mamba framework designed for spectral image reconstruction. GMSR-Net excels in achieving precise spectral reconstruction through the adept capture of long-range information utilizing a global receptive field. Notably, GMSR-Net strikes a favorable accuracy-efficiency trade-off, outperforming existing well-established methodologies. Compared with counterparts, GMSR-Net achieves a remarkable reduction in parameters and FLOPS, achieving a reduction by 10 times and 20 times, respectively.
@@
  • Extensive experiments on three spectral image reconstruction datasets show that the proposed GMSR-Net outperforms previous approaches in terms of efficiency and accuracy, providing a robust and promising lightweight architecture modeling solution for spectral reconstruction tasks.
Figure 2: Overview of GMSR-Net. (a) GMSR-Net.(b) Gradient Mamba. (c) Mamba Block. (d) Spatial Gradient Attention. (e) Spectral Gradient Attention.
@@ II Related Work
II-A Spectral Reconstruction from RGB.
Spectral reconstruction (SR) is an important technique aimed at reconstructing high spectral resolution HSIs solely from RGB images. It effectively addresses common challenges in hyperspectral imaging, such as high acquisition cost and low spatial resolution. In recent years, SR techniques have been rapidly developed from linear interpolation to sparse recovery. We summarize the SR algorithms into two categories: a priori-based methods and deep learning-based methods.
Most prior-based methods first perform statistical analysis on the data to obtain the intrinsic properties and features of the image, which are then utilized for SR. Since HSI exists in the form of a data cube, prior knowledge of it typically includes spatial structural similarity, sparsity, and correlations between spectra. Considering the sparsity of HSIs, Arad et al [18] proposed a sparse coding-based method to reconstruct HSIs from RGB images. First, compute the spectral dictionary, then project it into the RGB space, and finally reconstruct the HSI. The sparse representation based on a sparse dictionary accomplishes the task of reconstructing individual pixels without considering the spatial correlation, which limits the quality of the reconstructed image. To solve the above problem, Geng et al [19] proposed a spatially constrained dictionary learning SR method. By spatially introducing contextual information in the sparse representation, the spatial context of pixels is utilized to improve the reconstruction performance. However, HSI requires high spectral information, and the above method’s retention of spectral information needs to be enhanced. Given that the spectral signal is usually a relatively smooth function of wavelength, the method [20] employs a Gaussian process to model the spectral signal of each material and combines it with an RGB image in order to obtain a reconstructed HSI. Although this method improves the accuracy of the reconstruction, its solution process is relatively complex.
@@ III-A Overall Architecture
As shown in Fig.2, (a) introduces the proposed GMSR-Net, which is composed of multiple Gradient Mamba (GM) Blocks cascaded together. GMSR-Net takes a three-channel RGB image as input and aims to reconstruct the corresponding HIS. For the input image $RGB\in\mathbb{R}^{H\times W\times 3}$, firstly, channel upsampling is performed through a $1\times 1$ convolution to obtain the feature map $F_{1}\in\mathbb{R}^{H\times W\times\frac{C}{n}}$. Next, $F_{1}$ is input into the GM block (Fig.2 (b)). The GM block consists of three branches, which are the VSS (Visual State Space) block (Fig.2 (c)), Spatial Gradient Attention block (Fig.2 (d)), and Spectral Gradient Attention block (Fig.2 (e)). Wherein, the VSS block models long-range dependency relationships, while Spatial Gradient Attention and Spectral Gradient Attention sharpen structural information from all dimensions. Three feature maps $x_{1}$, $x_{2}$, and $x_{3}$ are obtained through the GM block and concatenated to get $y$. Finally, $y$ is channel downsampled and added with $x$ through skip connections to preserve the details and semantic information in the original input data, resulting in feature $z$. GMSR-Net gradually reconstructs spectral information by repeating the above process and ultimately obtains HIS.
@@ III-B VSS block
$\begin{array}[]{l}x^{\prime}=LN[SS2D(SiLU((DWConv(Lin(x)))))],\\ x^{\prime\prime}=SiLU(Lin(x)),\\ {x_{1}}=Lin(x^{\prime}\odot x^{\prime\prime}),\end{array}$ (1)
@@
SS2D. SS2D, the core of the VSS module, draws on the notion of a ”selectively scanned state-space sequence model” (S6), which was originally applied to natural language processing. Unlike traditional attention mechanisms, S6 enables each element in a sequence to interact with any previously scanned sample, thereby reducing the complexity of the attention mechanism from quadratic to linear. However, due to the non-causal nature of visual data, directly applying the S6 method to chunked and flattened visual images can lead to a loss of global receptive fields. To address this issue, we followed the approach outlined in [36] and adopted a ”four-direction” scanning strategy. As depicted in Fig.3, to better utilize 2D spatial information, we introduced the 2D-Selective-Scan (SS2D) operation, which flattens the 2D image features into a 1D sequence and scans along four different directions. This allows us to capture the long-range dependencies of each sequence. Finally, all sequences are merged using a summation method and reshaped to restore the 2D structure.
@@ III-C Spatial Gradient Attention
Although flattening spatial data into 1-D tokens can capture long-range dependencies, it disrupts the local 2-D dependencies of the image, thereby weakening the model’s ability to accurately interpret spatial relationships. While SS2D partially addresses this issue by scanning images horizontally and vertically, it still struggles to maintain the proximity of original adjacent tokens during the scanning sequence, resulting in the destruction of local details such as edges and textures. To solve the above problems, we propose spatial gradient attention to further deepen the contours in HIS and increase the clarity of HSI.
@@
Next, we perform channel-wise global max pooling and global average pooling operations on ${{\rm{G}}_{{\rm{spa}}}}$, obtaining two results, ${{\rm{G}}_{{\rm{spa}}}}_{{\rm{\_max}}}$ and ${{\rm{G}}_{{\rm{spa}}}}_{{\rm{\_avg}}}$, respectively. Then, these two results are concatenated along the channel dimension. Subsequently, a convolution operation is applied to reduce the number of channels to 1. Finally, spatial gradient attention feature ${\rm{GM}}{{\rm{A}}_{{\rm{spa}}}}$ is generated through the sigmoid function. Ultimately, ${\rm{GM}}{{\rm{A}}_{{\rm{spa}}}}$ is element-wise multiplied with the input features of the module, yielding spatially enhanced output features $F_{2}$.

### v2

@@ Abstract
Mainstream approaches to spectral reconstruction (SR) primarily focus on designing Convolution- and Transformer-based architectures. However, CNN methods often face challenges in handling long-range dependencies, whereas Transformers are constrained by computational efficiency limitations. Therefore, constructing a efficient SR network while ensuring the quality of reconstructed hyperspectral images (HSIs) has become a major challenge. Recent breakthroughs in the state-space model (e.g., Mamba) have attracted significant attention from natural language processing to vision tasks due to its near-linear computational efficiency and superior performance, prompting our investigation into its potential for SR problems. To this end, we propose the Gradient-guided Mamba for Spectral Reconstruction from RGB Images, dubbed GMSR-Net. GMSR-Net is a lightweight model characterized by a global receptive field and linear computational complexity. Its core comprises multiple stacked Gradient Mamba (GM) blocks, each featuring a tri-branch structure. Building upon the efficient global feature representation from the Mamba, we further innovatively propose spatial gradient attention and spectral gradient attention to guide the reconstruction of spatial and spectral cues. GMSR-Net demonstrates a significant accuracy-efficiency trade-off, achieving state-of-the-art performance while markedly reducing the number of parameters and computational burdens. Compared to existing approaches, GMSR-Net slashes parameters and FLOPs by substantial margins of 10 times and 20 times, respectively. Code is available at https://github.com/wxy11-27/GMSR.
@@
Lightweight Spectral Reconstruction, Mamba, State Space Model, Gradient Attention.
@@ I Introduction
Hyperspectral imaging is a technique utilized for acquiring images by capturing spectral information of an object across multiple contiguous wavelengths. Hyperspectral images (HSIs) can provide spectral information beyond the visible range of the human eye, thus making it possible to analyze and identify the properties, composition, and state of objects in greater detail. The fine spectral information of HSIs can substantially enhance the accuracy of feature classification and identification and is widely used for target identification [1], environmental monitoring [2, 3], and resource investigation [4]. However, hyperspectral imaging devices are often expensive, leading to high data acquisition costs, which limits their widespread use in certain fields. Additionally, due to the inherent physical properties of hyperspectral imaging sensors, HSIs always suffer from complex noise, low signal-to-noise ratio, and low spatial resolution, greatly restricting the accuracy improvement of high-level remote sensing tasks. Snapshot Compressive Imaging (SCI) systems and computational reconstruction algorithms [5, 6, 7] are effective means for reconstructing three-dimensional hyperspectral image (HSI) cubes. However, these methods still rely on expensive hardware equipment. As a cost-effective alternative, spectral super-resolution or spectral reconstruction (SR) techniques have attracted attention. SR techniques use color information and known spectral features in RGB images to estimate the spectral response curve of each pixel point to generate HSIs. The current SR methods can roughly be divided into sparse dictionary-learning [8, 9] and deep learning-based methods [10, 11].
@@
Figure 1: PSNR $v.s.$ Params $v.s.$ FLOPs comparisons with existing spectral reconstruction approaches. For an intuitive analysis, FLOPs and PSNR are represented by the horizontal axis and vertical axis, and the circle radius indicates Params. The proposed Gradient-Guided Mamba (GMSR-Net) outperforms counterparts with dramatically lower FLOPs and Params requirements.
The early SR research primarily focused on establishing sparse coding or relatively shallow learning models within specific spectral bands. However, due to their limited expressive power, these methods could only be effective within constrained image domains. In recent years, advancements in deep learning have significantly propelled the development of SR, overcoming many limitations and achieving significant progress. Convolutional neural networks (CNNs) have become the mainstream solution for SR due to their powerful representation and fitting capabilities. Current CNN-based methods treat SR as a regression problem, achieving it by establishing a mapping between HSIs and RGB images. HRNet [10] uses an encoder-decoder architecture and adds residual global blocks to extend the perceptual range. It won the track 2-Real World in the NTIRE 2020 Challenge Reconstructing Spectra from RGB Images. RepCPSI [12] significantly improves SR accuracy and reduces computational overhead by developing a lightweight coordinate-preserving proximity spectral-aware attention (CPSA) block. While CNN-based SR methods have made impressive progress in performance, CNNs are primarily adept at capturing local features. In the SR task, RGB images provide limited color information, while HSIs contain richer spectral information, which makes it necessary for the model to be able to efficiently map the information in the RGB image into the high-dimensional spectral space. Therefore, the model needs to be able to capture the long-range dependencies between different locations in the input image to ensure spatial consistency and spectral continuity when generating HSIs. In this regard, CNNs appear to be relatively limited.
@@
The Transformer [13] model was originally designed for natural language processing (NLP) tasks, but has since been successfully applied to computer vision. Its advantage lies in its ability to handle long-range dependencies and focus on global information. Cai et al. [6] first proposed an end-to-end Transformer-based framework MST, which aims to reconstruct HSI through compressed measurements. They proposed spectral-wise multi-head self-attention (S-MSA) to capture inter-spectra similarity and dependencies. Building on this, Lin et al. [14] embedded the sparsity of HSI into the Transformer model and developed a fine learning scheme for spectral synthesis imaging. However, since the Transformer needs to compute the attention mechanism between each image patch, the complexity of computation increases exponentially as the resolution increases. Therefore, Transformer models typically have a large number of parameters, which can lead to challenges in training and deployment in resource-constrained situations. Constrained by computational resources, it is usually necessary to divide the image into smaller patches. This pre-processing step inevitably results in each patch contains only a portion of the object, thus providing limited contextual information, which damages the accuracy of reconstruction.
@@
Recently, a novel architecture named Mamba [15] has attracted widespread attention. Compared to traditional Transformers, Mamba demonstrates greater efficiency in handling long sequences while maintaining training and inference effectiveness. Scholars have begun to introduce Mamba into the vision fields, including remote sensing image processing, achieving innovative progress in tasks such as image classification [16] and semantic segmentation [17]. However, the aforementioned tasks emphasize high-level information in images, such as objects, scenes, and semantic labels. The potential of Mamba to reconstruct low-level information such as colors, textures, and edges remains largely unexplored and warrants further investigation. This observation also motivates us to explore the capabilities of Mamba in modeling within HSI SR networks.
@@
Mamba requires scanning the image row-by-row or column-by-column, which leads to the separation of tokens that are otherwise adjacent in 2D space, thus increasing the difficulty of feature extraction in localized regions. However, in HSI, local details such as edges and textures are crucial for subsequent analysis and understanding. The gradient map is a great means to effectively reflect the edge and texture information in an image. Currently, the extraction and utilization of image gradients has been widely recognized as a key technique in several fields such as target detection [18] and image segmentation [19]. For example, Canny [18] has become a classical edge detection method based on gradient information by utilizing the gradient information of an image at multiple scales. The SPSR [20] employs an additional gradient branching network to generate a high-resolution gradient map as an additional structural prior, thus alleviating the geometric deformation problem. Distinguished from natural images, HSIs contain not only spatial structural information but also rich spectral information. The gradient map of the spectral dimension can reveal patterns and trends between different wavelengths, assisting models in better understanding and capturing variations among spectral features. This helps improve the model’s comprehension and analytical capabilities regarding image content. Currently, the utilization of gradient information in the spectral dimension of HSIs remains relatively limited, and there is still room for further improvement in spectral feature extraction.
@@
Overall, the current SR task faces the following challenges: CNN- and Transformer-based methods struggle to balance accuracy and efficiency, and geometric distortions often occur during the reconstruction process. In order to more efficiently model the long-range dependencies in HSIs with limited computational resources and mitigate the spatial structure deformation, as well as to better understand and capture the variations and correlations between spectrals, we introduce the state-space model in the spectral reconstruction task for the first time and propose GMSR-Net. GMSR-Net is an efficient SR model with a global receptive field and linear complexity, consisting of multiple Gradient-Guided Mamba (GM) blocks. Each GM block utilizes a three-branch structure. In addition to facilitating efficient global feature extraction through the Mamba block, we propose two innovative attention mechanisms: spatial gradient attention and spectral gradient attention. These mechanisms are designed to meticulously preserve structural information across both spatial and spectral dimensions, enabling SSM to be compatible with SR task. It is worth mentioning that GMSR-Net successfully balances accuracy and computational efficiency. As shown in Fig. 1, GMSR-Net achieves state-of-the-art (SOTA) performance while significantly reducing parameters and computational overhead. The main contributions of this paper are summarized as follows:
@@
  • We propose GMSR-Net, the first Mamba framework designed for SR task. GMSR-Net excels in achieving precise SR through the adept capture of long-range information utilizing a global receptive field. Notably, GMSR-Net strikes a favorable accuracy-efficiency trade-off, outperforming existing well-established methodologies. Compared with counterparts, GMSR-Net achieves a remarkable reduction in parameters and FLOPs, achieving a reduction by 10 times and 20 times, respectively.
@@
  • Extensive experiments on three spectral image reconstruction datasets show that the proposed GMSR-Net outperforms previous approaches in terms of efficiency and accuracy, providing a robust and promising lightweight architecture modeling solution for SR tasks.
Figure 2: Overview of GMSR-Net. (a) GMSR-Net.(b) Gradient Mamba. (c) VSS Block. (d) Spatial Gradient Attention. (e) Spectral Gradient Attention.
@@ II Related Work
II-A Spectral Reconstruction from RGB Images.
Spectral reconstruction (SR) is an important technique aimed at reconstructing high spectral resolution HSIs only from RGB images. It effectively addresses common challenges in hyperspectral imaging, such as high acquisition cost and low spatial resolution. In recent years, SR techniques have been rapidly developed from linear interpolation to sparse recovery. We summarize the SR algorithms into two categories: a priori-based methods and deep learning-based methods.
Most prior-based methods first perform statistical analysis on the data to obtain the intrinsic properties and features of the image, which are then utilized for SR. Since HSI exists in the form of a data cube, prior knowledge of it typically includes spatial structural similarity, sparsity, and correlations between spectra. Arad et al [21] proposed a sparse coding-based method to reconstruct HSIs from RGB images. The sparse representation based on a sparse dictionary accomplishes the task of reconstructing individual pixels without considering the spatial correlation, which limits the quality of the reconstructed image. To solve the above problem, Geng et al [22] proposed a spatially constrained dictionary learning SR method. By spatially introducing contextual information in the sparse representation, the spatial context of pixels is utilized to improve the reconstruction performance. However, HSI requires high spectral information, and the above method’s retention of spectral information needs to be enhanced. Given that the spectral signal is usually a relatively smooth function of wavelength, the method [23] employs a Gaussian process to model the spectral signal of each material and combines it with an RGB image in order to obtain a reconstructed HSI. Although this method improves the accuracy of the reconstruction, its solution process is relatively complex.
@@ III-A Overall Architecture
As shown in Fig. 2, (a) introduces the proposed GMSR-Net, which is composed of multiple Gradient Mamba (GM) Blocks cascaded together. GMSR-Net takes a three-channel RGB image as input and aims to reconstruct the corresponding HSI. For the input image $RGB\in\mathbb{R}^{H\times W\times 3}$, firstly, channel upsampling is performed through a $1\times 1$ convolution to obtain the feature map $F_{1}\in\mathbb{R}^{H\times W\times\frac{C}{n}}$. Next, $F_{1}$ is input into the GM block (Fig. 2 (b)). The GM block consists of three branches, which are the VSS (Visual State Space) block (Fig. 2 (c)), Spatial Gradient Attention block (Fig. 2 (d)), and Spectral Gradient Attention block (Fig. 2 (e)). Wherein, the VSS block models long-range dependency relationships, while Spatial Gradient Attention and Spectral Gradient Attention sharpen structural information from all dimensions. Three feature maps $x_{1}$, $x_{2}$, and $x_{3}$ are obtained through the GM block and concatenated to get $y$. Finally, $y$ is channel downsampled and added with $x$ through skip connections to preserve the details and semantic information in the original input data, resulting in feature $z$. GMSR-Net incrementally reconstructs spectral information through the above iterative processes, ultimately yielding a high-quality HSI.
@@ III-B VSS block
$\begin{array}[]{l}x^{\prime}=LN[SS2D(SiLU((DWConv(Lin(x)))))]\\ x^{\prime\prime}=SiLU(Lin(x))\\ {x_{1}}=Lin(x^{\prime}\odot x^{\prime\prime})\end{array}$ (1)
@@
SS2D. SS2D, the core of the VSS block, draws on the notion of a ”selectively scanned state-space sequence model” (S6), which was originally applied to NLP. Unlike traditional attention mechanisms, S6 enables each element in a sequence to interact with any previously scanned sample, thereby reducing the complexity of the attention mechanism from quadratic to linear. However, the non-causal nature of visual data poses a challenge when applying the S6 method directly to chunked and flattened visual images, potentially resulting in a loss of global receptive fields. To tackle this issue, we followed the approach outlined in [38] and implemented a “four-direction” scanning strategy. As illustrated in Fig. 3, in order to effectively utilize 2D spatial information, we introduced the 2D-Selective-Scan (SS2D) operation. This operation flattens the 2D image into a 1D sequences and scans along four different directions, enabling the capture of long-range dependencies within each sequence. Subsequently, all sequences are merged by summation and reshaped to the original 2D structure.
@@ III-C Spatial Gradient Attention
Although flattening spatial data into 1D tokens endowing capturing long-range dependencies, it disrupts the local 2D dependencies of the image, thereby weakening the model’s ability to accurately interpret spatial relationships. While SS2D partially addresses this issue by scanning images horizontally and vertically, it still struggles to maintain the proximity of original adjacent tokens during the scanning sequence, resulting in the destruction of local details such as edges and textures. To solve the above problems, we introduce spatial gradient attention to further enhance the local cues (e.g., contours) in HSIs and improve their clarity.
@@
Next, we perform channel-wise global max pooling and global average pooling operations on ${{\rm{G}}_{{\rm{spa}}}}$, obtaining two results, ${{\rm{G}}_{{\rm{spa}}}}_{{\rm{\_max}}}$ and ${{\rm{G}}_{{\rm{spa}}}}_{{\rm{\_avg}}}$, respectively. Then, these two results are concatenated along the channel dimension. Subsequently, a convolution operation is applied to reduce the number of channels to 1. Finally, spatial gradient attention feature ${\rm{GM}}{{\rm{A}}_{{\rm{spa}}}}$ is generated through the sigmoid function. Ultimately, ${\rm{GM}}{{\rm{A}}_{{\rm{spa}}}}$ is element-wise multiplied with the input features of the module, yielding spatially enhanced output feature $F_{2}$.